# Directional anomaly detection

### v1

@@ Abstract
Semi-supervised anomaly detection is based on the principle that potential anomalies are those records that look different from normal training data. However, in some cases we are specifically interested in anomalies that correspond to high attribute values (or low, but not both). We present two asymmetrical distance measures that take this directionality into account: ramp distance and signed distance. Through experiments on synthetic and real-life datasets we show that ramp distance performs as well or better than the absolute distance traditionally used in anomaly detection. While signed distance also performs well on synthetic data, it performs substantially poorer on real-life datasets. We argue that this reflects the fact that in practice, good scores on some attributes should not be allowed to compensate for bad scores on others.
@@
nomaly detection, average localised proximity, distance measures, nearest neighbour distance, one-class classification, risk factors
@@ 1 Introduction
A defining characteristic of semi-supervised anomaly detection (also known as one-class classification [22]) is that the training set only contains normal data, but that we still want to create a model that can distinguish between normal and anomalous data. Potential reasons for working in this setting include not wanting to assume that the available examples of anomalies are representative of all anomalies, and wanting to flag all data that is different from normal data as a potential anomaly. Examples of application domains of anomality detection include failure detection in industrial settings and screening patients in healthcare.
@@
In the present paper, we are specifically concerned with quantitative tabular data, where the training set consists of a representative sample of a certain normal target class. In this context, all semi-supervised anomaly detection algorithms are necessarily based on the principle that anomality increases with distance away from the normal training data in the feature space. However, there are circumstances in which we possess additional helpful domain knowledge. In this paper, we consider how best to deal with the knowledge that only relatively high values of an attribute should be indicative of anomality, not relatively low values (or vice-versa). For instance, certain attributes may correspond to higher strain of a machine, or encode known risk factors for patients. In such cases, we may want to decide that we are only interested in detecting anomalies that express high strain, or high risk factors, i.e., that we want to detect machine failure and at-risk patients, rather than underutilisation and exceptionally healthy patients.
To this end, we propose adaptations of two effective anomaly detection algorithms, Nearest Neighbour Distance (NND) and Average Localised Proximity (ALP), that take this directionality into account.
@@
Other attempts to take domain knowledge into account include contextual anomaly detection, where anomality is conditioned on one or more contextual attributes [19, 14], and fair anomaly detection, where the goal is to ensure that anomalies follow the same distribution over one or more sensitive attributes as normal records [16, 26, 18]. However, to the best of our knowledge, this is the first work to investigate directional anomaly detection.
2 Background: NND and ALP
@@
Semi-supervised anomaly detection algorithms, also known as one-class classifiers or data descriptors, take a training set $X$ consisting of normal data and learn a model that assigns an anomaly score to new instances. In the present paper, we will use two such algorithms: Nearest Neighbour Distance (NND) and Average Localised Proximity (ALP), both of which are directly based on distance measurements in the feature space.
@@ 2.1 NND
NND is one of the simplest anomaly detection algorithms and goes back to at least [10]. In NND, the anomaly score of a test record corresponds to the distance to its $k$th nearest neighbour in the training set. Despite its simplicity, NND performs surprisingly well across diverse datasets [12], making it an attractive baseline. In [11], a slight modification was proposed, taking the linearly weighted average of the first $k$ nearest neighbour distances. Thus, the anomality score of a test record $y$ becomes
@@ 2.2 ALP
for a choice of $k\geq 1$ and weight vector $w$, where the weighted maximum $\operatorname*{\mathnormal{w}\,max\,}$ is defined by
@@
for any collection $X$, where $X^{(i)}$ is the $i$th largest element of $X$; and where $\operatorname{lp}_{i}(y)$ is the $i$th localised proximity of $y$ to the training set, defined by
@@
where $d_{i}(y)$ is the $i$th nearest neighbour distance of $y$ in the training set according to some distance measure $d$, and $D_{i}(y)$ is the average $i$th nearest neighbour distance in the training set local to $y$, defined as
@@
2.3 Hyperparameters
For both NND and ALP, the values $k$ and $l$, the various weights and the distance measure are hyperparameters, which can in principle be tuned when there are anomalities available for validation purposes [13]. In general, however, we have to use sensible default values.
A typical choice for the distance $d$ is a form of the Minkowski $p$-distance:
$d(y,x)=\left(\sum_{j\leq m}\left\lvert y_{j}-x_{j}\right\rvert^{p}\right)^{\frac{1}{p}},$ (6)
for some value $p\geq 0$. The two most frequently used measures are $p=2$, which is Euclidean distance, and $p=1$, which is Boscovich distance, also known as city-block, Manhattan or rectilinear distance. For NND, Boscovich distance generally outperforms Euclidean distance [12, 11], while for ALP this is the standard choice anyway [12]. For both weighted NND and ALP, the default choice is to use linearly descending weights everywhere [12, 11]. For unweighted NND, $k=1$ is optimal [12], but for weighted NND slightly higher values like $k=8$ increase general performance [11]. For ALP, $k=5.5\log n$ and $l=6\log n$ have been established as optimal default values [12], where $n$ is the number of training records.
While not typically seen as a hyperparameter, both NND and ALP are also dependent on the relative scale of the attributes.
3 Directional anomaly detection
@@
As discussed in the introduction, we want to find a way to take into account directional attributes, for which we should only interpret extreme values in one direction as anomalous. Without loss of generality, we will assume that high values correspond with anomality — any attributes for which the opposite applies can be transformed with a sign change. We will use the following potentially asymmetric distance measure: [^1] Note that while this introduces ramp distance and signed distance as variants of Boscovich distance, it is also possible to substitute ramp distance in the calculation of Minkowski distance (6) with other values of $p$, and in particular for Euclidean distance. For signed distance, adapting Minkowski distance requires more work.
$d(y,x)=\sum_{j\leq m}d_{j}(y_{i}-x_{i}),$ (7)
and consider three variants of the per-attribute distance $d_{j}(y_{i}-x_{i})$, listed in Table 1. Absolute distance is the baseline. It corresponds to the existing Boscovich distance, i.e. Minkowski distance (6) with $p=1$. Ramp distance is the result of applying the ramp function to the difference $y_{j}-x_{j}$. The ramp function is also known as the rectifier function, or rectified linear unit (ReLU), which is a popular activation function in neural networks [4]. Signed distance is simply the identity function applied to $y_{j}-x_{j}$.
Table 1: Variant approaches to directional attributes.
Variant $d_{j}(y_{i}-x_{i})$
Absolute $\left\lvert y_{j}-x_{j}\right\rvert$
Ramp $\max(0,y_{j}-x_{j})$
Signed $y_{j}-x_{j}$
All three variants agree for positive values of $y_{j}-x_{j}$; they interpret this as positive evidence for the anomality of $y$. They differ with respect to negative values of $y_{j}-x_{j}$, i.e., they differ in how to account for test values that are lower than the reference value in the training set. Absolute distance counts such low values as positive evidence for the anomality of $y$ just as it does high values, thereby ignoring the directionality of the attribute $j$. In contrast, signed distance interprets such low values as negative evidence for anomality, which can offset high values in other attributes. Finally, ramp distance stakes out the middle ground by interpreting such values as a lack of evidence either way.
Next, we consider how these variant distances affect NND and ALP. We can simply substitute the ramp distance into the overall distance for all directional attributes, and calculate the nearest neighbour distances according to the resulting measure.
For signed distance, the situation is not so straightforward. Note first that when all attributes are directional and we use signed distance, the nearest neighbour of any test record $y$ is the training instance $x$ that minimises $\sum_{i}(y_{i}-x_{i})$. This is the training instance with the largest attribute value sum $\sum_{i}x_{i}$, and this does not depend on the test instance. Analogously, all test instances share the same subsequently nearest neighbours. Consequently, we can disregard nearest neighbours altogether, and simply compare the attribute sums of test records. Note that we still use the training set to determine the right scale of each attribute. Nevertheless, because we no longer need to perform any nearest neighbour queries we end up with a significant computational simplification.
Figure 1: A dataset with one directional and one adirectional attribute, one test record $y$ and two training records $x$ and $x^{\prime}$. See text for discussion.
What should we do when we have a mixture of adirectional and directional attributes? One option is to add the signed distance from the directional attributes and the absolute distance from the adirectional attributes together to obtain a mixed distance between every test and training record and use this to identify the nearest neighbour distances of every test record. However, this would lead to the paradoxical situation illustrated by Fig. 1. The nearest neighbour of $y$ is $x^{\prime}$, since its mixed distance $4-6=-2$ is smaller than the mixed distance $2+0=2$ between $y$ and $x$. However, in terms of absolute distance, $y$ is clearly more similar to $x$, even if we only consider the adirectional attribute, and so it seems nonsensical to calculate the anomality of $y$ on the basis of its distance to $x^{\prime}$.
Therefore, we propose that for NND modified with signed distance, the right order of things is to seperately calculate the risk score for the directional attributes and the nearest neighbour distance for the adirectional attributes and add these together to obtain a single anomality score. Because of its effect on nearest neighbour calculations, we will not attempt to use signed distance with ALP.
Figure 2: Mean AUROC obtained with NND for synthetically generated datasets with Gaussian and Bernoulli attributes. $a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
Figure 3: Mean AUROC obtained with ALP for synthetically generated datasets with Gaussian and Bernoulli attributes. $a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
4 Experiments on synthetic datasets
In order to evaluate our proposal, we construct a series of synthetic datasets, each containing 1000 normal training records and 200 test records — 100 normal and 100 anomalous. Each record consists of 10 attributes with randomly generated values. We will consider two types of attributes. First, we generate datasets with continuous attributes, where both normal and anomalous values follow a Gaussian distribution with standard deviation 1, centred on 0 and $a$ respectively. We let $a$ vary between 0 and 1 in steps of 0.1. Secondly, we generate datasets with binary attributes, where normal and anomalous values follow Bernoulli distributions with $p=0.5-0.5\cdot b$ and $p=0.5+0.5\cdot b$ respectively, for which we let $b$ vary between 0 and 0.5 in steps of 0.05. For each value of $a$ and each value of $b$ we generate 100 datasets.
@@ 4 Experiments on synthetic datasets
We evaluate anomaly detection performance using the area under the receiver operating characteristic (AUROC). This expresses the ability of an anomaly detector to separate anomalies from normal data. For this purpose, we monotonically transform the NND distance from (1) to an anomaly score in $[0,1]$ with:
@@
We do this for absolute distance, ramp distance and signed distance. The results for NND are displayed in Figure 2. For the Gaussian attributes, signed and ramp distance perform substantially better than absolute distance, for small and large $k$, in line with our directional hypothesis. Signed distance performs slightly better than ramp distance, but the difference is small. For Bernoulli attributes, the picture looks a bit different. At $k=1$, signed distance outperforms absolute distance, but ramp distance performs worse than absolute distance. At $k=8$, the gap between signed and absolute distance has tightened considerably, and ramp distance now sits in between. At $k=100$, signed and ramp distance have become indistinguishable in terms of performance, and absolute distance is only very slightly behind.
For ALP, the results for the default hyperparameter values $k=5.5\log 1000=38$ and $l=6\log 1000=41$ are displayed in Table 3. For both Gaussian and Bernoulli attributes, ramp distance outperforms absolute distance, although more clearly so for Gaussian attributes.
Taken together, these results suggest that when we know for a given attribute that anomalies have a higher mean value than normal data, we should indeed use signed or ramp distance rather than absolute distance. Between these two variants, signed distance appears to have the advantage with NND, but we will see in the next section that the situation is different for real-life datasets.

### v2

# Monotonic anomaly detection

@@ Abstract
Semi-supervised anomaly detection is based on the principle that potential anomalies are those records that look different from normal training data. However, in some cases we are specifically interested in anomalies that correspond to high attribute values (or low, but not both). We present two asymmetrical distance measures that take this monotonicity into account: ramp distance and signed distance. Through experiments on synthetic and real-life datasets, we show that ramp distance increases anomaly detection performance over the traditional absolute distance. While signed distance also performs well on synthetic data, it performs substantially poorer on real-life datasets. We argue that this is a consequence of the fact that when using signed distance, low values of certain attributes automatically compensate for high values of other attributes, such that anomaly detection is reduced to counting the total attribute value sum, which is too simplistic in practice.
@@
nomaly detection, distance measures, monotonicity, nearest neighbours, one-class classification, risk factors
@@ 1 Introduction
A defining characteristic of semi-supervised anomaly detection (also known as one-class classification [22]) is that the training set only contains normal records, but that we still want to create a model that can distinguish between normal records and anomalies. Potential reasons for working in this setting include not wanting to assume that the available examples of anomalies are representative of all anomalies, and wanting to flag all new records that deviate from normal data as potential anomalies. Examples of application domains of anomality detection include failure detection in industrial settings and screening patients in healthcare.
@@
In the present paper, we are specifically concerned with quantitative tabular data, where the training set consists of a representative sample of a certain normal target class. In this context, all semi-supervised anomaly detection algorithms are necessarily based on the principle that anomality increases with distance away from the normal training data in the feature space. However, there are circumstances in which we possess additional helpful domain knowledge.
In this paper, we consider how best to deal with the knowledge that only relatively high values of an attribute should be indicative of anomality, not relatively low values (or vice-versa). For instance, certain attributes may correspond to higher strain of a machine, or encode known risk factors for patients. In such cases, we may only be interested in detecting anomalies that express high strain, or high risk factors, i.e., machine failure and at-risk patients, rather than underutilisation and exceptionally healthy patients. We call this setting monotonic anomaly detection, because our domain knowledge places a monotonicity constraint on certain attributes and the prediction. Alternatively, we may think of this as directional anomaly detection, because the goal is to detect anomalies that lie in a certain direction from the normal data in the attribute space.
When applied to such a monotonic anomaly detection problem, an ordinary anomaly detection algorithm will needlessly flag records that represent underutilisation or exceptionally healthy patients. But the real problem is more fundamental. Existing anomaly detection algorithms will flag any record that deviates sufficiently from the normal records, due to a combination of relatively high values for some attributes and relatively low values for others. This is problematic if we believe that relatively low values should not be indicative of anomality.
Consequently, we propose that a correct use of monotonic domain knowledge requires that we adapt existing algorithms by modifying the contribution of individual attribute values to the anomaly score. Specifically, we have to ask what the contribution of a relatively low attribute value of a test record should be. Logically, there are three possibilities: relatively low attribute values can be interpreted as positive evidence of anomality, as negative evidence against anomality, and as neither positive nor negative evidence. The first of these three alternatives corresponds to ordinary anomality detection, because it makes no difference between relatively low and relatively high attribute values — both are treated as positive evidence of anomality. In contrast, the second and third option represent two alternative monotonic approaches, because they treat low and high attribute values differently.
In the present paper, we investigate how distance-based algorithms can be modified for monotonic anomaly detection. Specifically, we will consider Centre Distance (CD), Nearest Neighbour Distance (NND), Local Outlier Factor (LOF) and Average Localised Proximity (ALP). The reason for focusing on these algorithms is twofold. Firstly, they are all based on the distance between records, and we can make them monotonic by choosing an appropriate asymmetric distance measure. Secondly, these are some of the best-performing algorithms in the literature. In an experimental evaluation across 246 semi-supervised anomaly detection problems [12], ALP, NND and LOF placed first, third and fourth.
In Section 2, we briefly mention some related work, and in Section 3 we describe the CD, NND, LOF and ALP algorithms. Next, we present our proposed asymmetric distance measures that enable directional anomaly detection in Section 4. Finally, we describe our experimental setup in Section 5 and present the results in Section 6, before concluding in Section 7. We also include Appendix 0.A with an analysis of the hyperparameter choice $k$ for NND.
2 Related work
There exists a substantial literature on monotonic classification [3]. However, because approaches proposed in that context rely on a representative sample from each class in the training set, they cannot be applied to semi-supervised anomaly detection.
The closest related work to the present paper is the proposal on direction-related anomaly detection [25]. The problem setting in that work is a form of ordinal regression, where anomalies are the records with the highest response value, which are not represented in the training set. The proposed approach is to use an ordinal regression algorithm to learn a decision boundary beyond which lie the anomalies. Thus, the directionality in that setting is determined by the response values, rather than the independent attributes.
@@ 2 Related work
Other attempts to incorporate domain knowledge in anomaly detection include contextual anomaly detection, where anomality is conditioned on one or more contextual attributes [19, 14], and fair anomaly detection, where the goal is to ensure that anomalies follow the same distribution over one or more sensitive attributes as normal records [16, 27, 18].
3 Anomaly detection algorithms
@@
Semi-supervised anomaly detection algorithms, also known as one-class classifiers or data descriptors, take a training set $X$ consisting of normal data and learn a model that assigns an anomaly score to new instances. In the present paper, we will consider four such algorithms: Centre Distance (CD), Nearest Neighbour Distance (NND), Local Outlier Factor (LOF) and Average Localised Proximity (ALP).
3.1 CD
CD is arguably the simplest possible anomaly detection algorithm. It requires choosing a centre $c$, a single value such as the marginal mean or median that is representative of the whole training set. The anomaly score of a test record $y$ is its distance to $c$. Due to its simplicity, CD is a useful baseline.
@@ 3.2 NND
NND is also very simple, and goes back to at least [10]. The anomaly score of a test record corresponds to the distance to its $k$th nearest neighbour in the training set. Despite its simplicity, NND performs surprisingly well across diverse datasets [12], making it an attractive default choice for anomaly detection. In [11], a slight modification was proposed, taking the linearly weighted average of the first $k$ nearest neighbour distances. Thus, the anomality score of a test record $y$ becomes:
@@
3.3 LOF
The idea behind LOF is to estimate the density of the normal data at the location of a test record $y$, and to compare this to the density at the location of its nearest neighbours. The anomality score of $y$ corresponds inversely to this relative density. Concretely, it is calculated as follows:
$\frac{1}{k}\sum_{j\leq k}\frac{\operatorname{lrd}_{k}(\operatorname{NN}_{j}(y))}{\operatorname{lrd}_{k}(y)},$ (2)
for a choice of $k\geq 1$, and a distance measure $d$, where $\operatorname{NN}_{j}(y)$ is the $j$th nearest neighbour of $y$ in the training set, and $\operatorname{lrd}_{k}$ is the so-called local reachability density, which in turn corresponds inversely to the distance between a record and its neighbours:
$\operatorname{lrd}_{k}(y)=\frac{1}{\frac{1}{k}\sum_{i\leq k}\operatorname{rd}_{k}(y,\operatorname{NN}_{i}(y))},$ (3)
where $\operatorname{rd}_{k}$ is the so-called reachability distance:
$\operatorname{rd}_{k}(y,x)=\max(d(y,x),d_{k}(x)),$ (4)
where $d_{k}(x)$ is the $k$th nearest neighbour distance of $x$ in the training set. The motivation for using the asymmetric reachability distance rather than ordinary distance is that this results in a smoother density function.
LOF is a well-known and widely-used algorithm. Nevertheless, when evaluated across a wide range of real-life datasets, it failed to beat the simpler NND algorithm [12].
@@ 3.4 ALP
for a choice of $k\geq 1$ and weight vector $w$, where the weighted maximum $\operatorname*{\mathnormal{w}\,max\,}$ is defined by:
@@
for any collection $X$, where $X^{(i)}$ is the $i$th largest element of $X$; and where $\operatorname{lp}_{i}(y)$ is the $i$th localised proximity of $y$ to the training set, defined by:
@@
where $d_{i}(y)$ is the $i$th nearest neighbour distance of $y$ in the training set according to some distance measure $d$, and $D_{i}(y)$ is the average $i$th nearest neighbour distance in the training set local to $y$, defined as:
@@
In the large experiment on real-life datasets [12], ALP came out on top.
4 Monotonic anomaly detection
@@
As discussed in the introduction, we want to find a way to take into account monotonic attributes, for which we should only interpret extreme values in one direction as anomalous. Without loss of generality, we will assume that high values correspond to anomality — any attributes for which the opposite applies can be transformed with a sign change.
4.1 Asymmetric distance
The anomaly detection algorithms described in Section 3 are all based on a choice of distance measure $d$ between various records. For tabular data, a typical choice for the distance $d$ is a form of Minkowski $p$-distance:
$d(y,x)=\Big{(}\sum_{i\leq m}\lvert y_{i}-x_{i}\rvert^{p}\Big{)}^{\frac{1}{p}},$ (9)
for some $p\in(0,\infty)$. [^1] In addition, Minkowski $\infty$-distance, also known as Chebyshev distance, is the limit case $p\rightarrow\infty$, which resolves to $d(y,x)=\max_{i\leq m}\left\lvert y_{i}-x_{i}\right\rvert$. In particular, the two most frequently used choices are 1-distance and 2-distance.2221-distance is also known as Boscovich, city-block or Manhattan distance. 2-distance is also known as Euclidean distance.
Minkowski $p$-distance is a function of the per-attribute differences $y_{i}-x_{i}$. In the anomaly detection algorithms described in Section 3, the distance is applied to test records $y$ and training records $x$. When $y_{i}$ is larger than $x_{i}$, this contributes positive evidence to the anomality of $y$. And because the Minkowski $p$-distance takes the absolute value of the per-attribute differences, values of $y_{i}$ that are smaller than $x_{i}$ likewise contribute positive evidence for the anomality of $y$. However, when we are dealing with a monotonic attribute, small values of $y_{i}$ are not an indication of anomality.
Logically, there are two alternative ways that values of $y_{i}$ smaller than $x_{i}$ can be interpreted. By not taking the absolute value, but simply using the signed difference $y_{i}-x_{i}$, we interpret small values of $y_{i}$ as negative evidence against anomality. And by applying the ramp function [^3] This is also known as the rectifier function, or rectified linear unit (ReLU), and is a popular activation function in neural networks [5]. $\max(0,y_{i}-x_{i})$, we do not interpret small values of $y_{i}$ as providing evidence either for or against anomality.
In general, a dataset may contain a mixture of monotonic and ordinary attributes. Therefore, we propose the following generalised, potentially asymmetric form of Minkowski $p$-distance:
[…]
We will use absolute distance for all ordinary attributes, and compare the use of absolute, ramp and signed distance for monotonic attributes. Using absolute distance for monotonic attributes is the baseline, since this corresponds to traditional anomaly detection.
Table 1: Variant approaches to monotonic attributes.
Variant $\delta_{i}(y_{i}-x_{i})$ Interpretation of negative values
Absolute $\left\lvert y_{i}-x_{i}\right\rvert$ Evidence for anomality
Ramp $\max(0,y_{i}-x_{i})$ No evidence either way
Signed $y_{i}-x_{i}$ Evidence against anomality
4.2 Anomaly detection with signed distance
Whereas ramp distance can be straightforwardly substituted into the definitions of the anomaly detection algorithms described in Section 3, the use of signed distance leads to complications.
To simplify the analysis, we will use 1-distance, and consider a dataset that consists entirely of monotonic attributes, which is the case for the majority of datasets in our experiments.
For CD, we find that when we use signed distance, the choice of centre $c\in\mathbb{R}^{m}$ does not actually matter. By definition, the anomaly score of a test record $y\in\mathbb{R}^{m}$ is $\sum_{i\leq m}(y_{i}-c_{i})$. Therefore, a change in centre from $c$ to $c^{\prime}$ simply induces a shift of all anomaly scores by a constant $\sum_{i\leq m}(c_{i}-c^{\prime}_{i})$. Effectively, the anomaly score of $y$ is simply the sum of its attribute values, i.e. the total risk of a record. [^4] This reduces the role of the training set to scaling, i.e. to determine the relative weight of each risk factor.
Furthermore, the nearest neighbour of $y$ is by definition the training record $x$ that minimises $\sum_{i\leq m}(y_{i}-x_{i})$, which is the training record $x$ with the highest attribute value sum $\sum_{i\leq m}x_{i}$. This is independent of $y$, meaning that we lose the local character of nearest neighbour algorithms. In particular, NND with signed distance is equivalent to CD, since we saw above that the choice of centre does not fundamentally matter for the anomaly score of CD, and we can choose $x$ to be this centre.
In light of this, we can only evaluate signed distance for CD.
5 Experimental setup
We will conduct a series of experiments on synthetic and real-life datasets from the UCI repository [4] to evaluate whether the approach proposed in this paper leads to better anomaly detection performance on datasets with monotonic attributes. Concretely, we will compare the use of absolute distance (ordinary anomaly detection) with ramp distance (monotonic anomaly detection), as well as signed distance for the CD algorithm.
5.1 Evaluation
@@
We evaluate anomaly detection performance using the area under the receiver operating characteristic (AUROC). This expresses the ability of an anomaly detector to separate anomalies from normal data. For this purpose, we monotonically transform the NND score from (1) and the LOF score (2) to an anomaly score in $[0,1]$ with:
$a\longmapsto\frac{a}{\left\lvert a\right\rvert+1}.$ (11)
Because the CD score can be negative when using signed distance, we instead use the following function to transform it to an anomaly score in $[0,1]$:
@@
5.2 Synthetic datasets
The synthetic datasets contain 1000 normal training records and 200 test records — 100 normal and 100 anomalous. Each record consists of 10 attributes with randomly generated values. We will consider two types of attributes. First, we generate datasets with continuous attributes, where both normal and anomalous values follow a Gaussian distribution with standard deviation 1, centred on 0 and $a$ respectively. We let $a$ vary between 0 and 1 in steps of 0.1. Second, we generate datasets with binary attributes, where normal and anomalous values follow Bernoulli distributions with $p=0.5-0.5\cdot b$ and $p=0.5+0.5\cdot b$ respectively, for which we let $b$ vary between 0 and 0.5 in steps of 0.05. For each value of $a$ and each value of $b$ we generate 100 datasets, and calculate the mean AUROC across these datasets.